OpenStudy (anonymous):

Solve the equation:

OpenStudy (anonymous):

\[4\sin ^{2}x = 3\]

OpenStudy (anonymous):

Use the identity: \[\sin ^{2}x -1 = -\cos(2x)\] 4cos(2x) = 2 cos(2x) = 0.5 \[\cos^{-1} (0.5) = 2x\] x = \[ \frac{\cos^{-1}(0.5) }{ 2 }\]

OpenStudy (anonymous):

oops i forgot the negative

OpenStudy (anonymous):

piglet can u helpme onmy question plz?